# On Sensitivity of Learning with Limited Labelled Data to the Effects of Randomness: Impact of Interactions and Systematic Choices

### v1

@@ B.1 Formal Description of the Method
Afterwards, a model is trained and evaluated multiple ($N$) times, each time with a different configuration $n$ from $\mathbb{IFC}_{i}$ ($n\in\mathbb{IFC}_{i}$), while at the same time keeping the configuration of all the mitigated randomness factors fixed. To achieve this, a single configuration $m$ is randomly sampled from $\mathbb{MFC}_{i}$ ($m\in\mathbb{MFC}_{i}$) and used for each run. The model performance of each run ($r_{m,n}$) is determined. This training and evaluation constitutes a single investigation run, where we observe the changes introduced by the investigated randomness factor.
The standard deviation $p\_std_{m}$ of model performance across these runs ($p\_std_{m}=std(r_{m,*})$) represents the effects of the investigated randomness factor. As it is determined using a fixed configuration $m$ of the mitigated randomness factors, these effects are still affected by the interactions between the factors. Therefore, we call this standard deviation $p\_std_{m}$ as partial standard deviation.
To remove the effects of other randomness factors, the single investigation run is performed multiple times, calculating partial standard deviation for each configuration $m$ from $\mathbb{MFC}_{i}$. The final adjusted effects of the investigated randomness factor, denoted also as contributed standard deviation $c\_std$, are calculated as the average over the partial standard deviations ($c\_std=\overline{p\_std_{*}}$). Another possibility to remove the effects of other randomness factors is to use a specific mitigation strategy instead of the repeated training runs, such as using sample selection strategy to select the samples for the investigation runs instead of randomly. The method can be easily extended to use such strategy for all the mitigated factors or just their subset by simply changing how the randomness factor configuration is selected.
To assess the importance of the investigated randomness factor, the contributed standard deviation $c\_std$ is compared with two additional values: 1) mitigated standard deviation $m\_std$; and 2) golden model standard deviation $gm\_std$.
The mitigated standard deviation represents the joint effects of all the non-investigated randomness factors. A partial mean (denoted as $p\_mean_{m}$) is first calculated for each single investigation run (i.e., at the same time as the partial standard deviation $p\_std_{m}$). This partial mean represents the expected average model performance for the given combination of configurations of the non-investigated randomness factors. The mitigated standard deviation $m\_std$ is calculated as the standard deviation across all partial means from investigation runs that use different configurations from $\mathbb{MFC}_{i}$ ($m\_std=std(p\_mean_{*})$).
The golden model standard deviation represents an objective estimate of the deviation in the model performance. To get this estimate, the same investigation strategy is performed, but without any investigated randomness factor. Instead, a golden model configuration set $\mathbb{GMC}$ ($|\mathbb{GMC}|=L$) is defined, as a cross product between the sets of all the randomness factor configurations:
@@
Similar to the investigation, a model is trained and evaluated for every configuration $g$ from $\mathbb{GMC}$ and the model performance $r_{g}$ is determined. The standard deviation across these runs represents the golden model standard deviation $gm\_std$. Another possibility to determine this value is to calculate the standard deviation across all the model performance results $r_{m,n}$ obtained during investigation. As all the combination of randomness factor configurations are considered during investigation, it is possible to simply calculate the golden model standard deviation as a single standard deviation across all the model performance results $r_{m,n}$ obtained during investigation.
The final importance score of the investigated randomness factor is calculated as:
$importance=(c\_std-m\_std)/gm\_std$ (3)
@@
This importance score measures how much more of the portion of the overall standard deviation the investigated factor contributed over all the non-investigated ones. Any factor with the importance score of more than 00 is considered to be significantly important, as it contributes the same amount of deviation as all the remaining factors. In addition, higher importance score indicates that the factor is more important. As such, the importance score can be used to order the different randomness factors based on their importance and help focus on the most important ones.
In real world scenario, an approximation of this method needs to be used. The randomness factor configuration sets $\mathbb{C}_{j}$ may be extremely large or even infinite, making the search over all the possible configurations infeasible. As an approximation, a randomly selected subsets $\mathbb{IFC}_{i}^{{}^{\prime}}$, $\mathbb{MFC}_{i}^{{}^{\prime}}$ and $\mathbb{GMC}^{{}^{\prime}}$ may be used instead. The sizes of these subsets, denoted $N$, $M$ and $L$ respectively, provide a trade-off between feasibility (or computation costs) of the investigation and precision (or reliability) of the investigation results (an ablation study of this trade-off is included in the Appendix C). To achieve reliable estimate of the effects of randomness factors and their importance, the sizes should adhere to conditions defined at the end of Section 3. However, it is important to note that in case of extending the method, the requirements for reliable estimate of the effects of randomness factors also change. To investigate multiple randomness factors at the same time, more investigation runs should be used (i.e., size $N$ should be increased), while keeping the number of mitigation runs similar (i.e., size $M$). In case when a more sophisticated mitigation strategies are used, the number of mitigation runs may be reduced (i.e., size $M$).
B.2 Additional Resources Describing the Proposed Method
@@
To allow for better understanding of our proposed investigation method, we provide further resources that describe how it works. First of all, we provide more informal description of the steps composed in Algorithm 1, along with references to individual lines in it. Informally our proposed method works in a following way:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    The golden model standard deviation is calculated by simply performing training and evaluation multiple times with differently fixed configuration of all randomness factors. (final loop; lines 13-17 in the Algorithm 1)
  6. 6.
    The importance score of the investigated factor is calculated as a fraction of the golden model standard deviation of the difference between contributed standard deviation and the mitigated standard deviation (to determine how much more the investigated factor contributes over all the mitigated ones). Any randomness factor with importance score over 00 is considered significantly important, as such factors contribute the same amount of deviation as the combination of all the remaining factors. (the final check; lines 18-21 in the Algorithm 1)
Second, we provide the visualisation of the method in a form of table. In essence, when investigating the specific randomness factor, while mitigating the effects of other randomness factors, we fill in such table as illustrated in Table 2. The columns represent the different configurations for the investigated factor. Observing how the performance changes across these columns, we can determine the effects of the randomness factors – aggregating across these columns we obtain the partial mean $p\_mean$ and partial standard deviation $p\_std$.

### v2

@@ 3 Investigation of Randomness while Taking Interactions into Consideration
Investigating effects.
At its core, the investigation of factor $i$ is done by observing how the performance changes across the different configurations the randomness factor can appear in. In a single investigation run, the training and evaluation of a model is repeated $N$ times ($N=|\mathbb{IFC}_{i}|$), each time with a different configurations $n$ of the factor $i$, while keeping the configurations of the remaining factors fixed to a randomly chosen configuration $m$ from $\mathbb{MFC}_{i}$. For each repeat, the model performance ($r_{m,n}$) is determined. The standard deviation $p\_std_{m}$ (called partial standard deviation) across these $N$ runs ($p\_std_{m}=std(r_{m,*})$) represents the effects of the investigated randomness factor that are still affected by the interactions.
Mitigating interactions.
To remove the effects of other randomness factors, the investigation run is repeated multiple ($M$) times each time with a different fixed configuration $m$ from $\mathbb{MFC}_{i}$. Each such repeat is called mitigation run and results in a separate partial standard deviation. After performing enough mitigation runs (i.e., searching through enough configurations $m$ of the non-investigated randomness factors), the partial standard deviations ($p\_std_{m}$) are averaged to produce the contributed standard deviation ($c\_std=\overline{p\_std_{*}}$), which represents the final adjusted effects of the investigated factor $i$ (i.e., it represents the deviation the investigated randomness factor contributes to the overall deviation in results).
Calculating importance score.
To assess the importance of the factor, the contributed standard deviation is compared with two additional values: 1) mitigated standard deviation ($m\_std$); and 2) golden model standard deviation ($gm\_std$). The mitigated standard deviation represents the joint effects of all the non-investigated randomness factors (i.e., standard deviation contributed by non-investigated factors). To obtain this value, a partial mean ($p\_mean_{m}$) is calculated for each investigation run, which represents the expected average model performance for the given combination of configurations of the non-investigated factors. The mitigated standard deviation is then calculated as the standard deviation across these partial means ($m\_std=std(p\_mean_{*})$).
The golden model standard deviation ($gm\_std$) represents an objective estimate of the deviation in the model performance. To get this estimate, a golden model configuration set $\mathbb{GMC}$ ($|\mathbb{GMC}|=L$) is defined, as a cross product between the sets of all the randomness factor configurations:
@@
Afterwards, a model is trained and evaluated $L$ times each time with different configuration $g$ from $\mathbb{GMC}$, the model performance $r_{g}$ is determined and the standard deviation across these runs represents the golden model standard deviation $gm\_std$.
The final importance score of the factor is defined as the portion of the golden model standard deviation the investigated factors contribute over the non-investigated ones ($importance=(c\_std-m\_std)/gm\_std$). Any randomness factor with an importance value over 00 is considered important, as it contributes the same amount of deviation as all the remaining factors combined. The size of the score determines the relative importance between the factors (e.g., factor with importance score of 0.6 is more important than one with score of 0.1) and can be used for further analysis and comparison across different factors, models, datasets and experimental settings (e.g., how the importance of specific factor changes if the number of samples is increased or a different dataset is used).
Choosing values for parameters $N$, $M$ and $L$.
The number of investigation runs ($N$) and the mitigation runs ($M$) provide a trade-off between the feasibility (or computation costs) of the investigation and the precision of the results (how well the effects are estimated and interactions mitigated). Below, we provide a set of heuristics to achieve a good trade-off (and provide full method for selecting the values of the parameters in Appendix B.3):
  1. 1.
    $N,M\gg 1$; $N$ and $M$ should cover a large enough number of factor configurations to sufficiently estimate the effects.
  2. 2.
    $M\geq N$; as the higher number of mitigation runs ($M$) leads to better mitigation of the interactions, increasing value of $M$ should be preferred over increasing the number of investigation runs ($N$).
  3. 3.
    $L=N*M$; to guarantee the importance score is calculated from distributions of the same sizes and characteristics, the number of runs in the golden model should be equal to the overall number of runs in the investigation.
Validation of the proposed method.
We evaluate the validity of the proposed method indirectly (as there is no ground-truth to compare against) using the following experiments: 1) comparing the method to two existing baselines (i.e., Random and Fixed investigation strategy) and evaluating the properties and benefits of our method, specifically the handling of interactions that may lead to underestimation or overestimation of the effects in specific cases, and the importance score that allows for more in-depth analysis and comparison across different experimental settings; 2) exploring the dependence of how well the effects are estimated and their interactions mitigated by our method to the number of investigation and mitigation runs, where we found that the results of our methods are stable already with a low number of runs (20 mitigation and 10 investigation runs); and 3) observing the consistency of the results and findings when applying the method to different settings (factors, approaches, datasets). The full description of the validation results is in Appendix E.
4 Experiments
Datasets.
The experiments are conducted on 7 text classification datasets composed of different tasks with different number of classes. We focus on 3 binary classification datasets from the GLUE benchmark Wang et al. (2018): SST2 Socher et al. (2013) for sentiment classification, CoLA Warstadt et al. (2019) for determining the grammatical acceptability of a sentence, and MRPC Dolan and Brockett (2005) for determining the semantic equivalence relationship between two sentences. In addition, we use 4 multi-class text datasets: AG News Zhang et al. (2015) for news classification, TREC Voorhees and Tice (2000) for question classification, DB-Pedia Lehmann et al. (2015) for topic classification and SNIPS Coucke et al. (2018) for intent classification.
Approaches.
The main focus of the investigation is on the in-context learning using the Flan-T5 Chung et al. (2022) base, LLaMA-2 Touvron et al. (2023) 13B instruction optimised model, Mistral-7B Jiang et al. (2023) and Zephyr-7B Tunstall et al. (2023). In addition, we also focus on fine-tuning, using the BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) base models. Finally, we also investigate the meta-learning approaches MAML Finn et al. (2017), Reptile Nichol et al. (2018) and the Prototypical Networks Snell et al. (2017), but only on the binary datasets.
Randomness Factors.
Method Setup.
For each randomness factor, the number of the investigation runs ($N$) is set to 10, the number of mitigation runs ($M$) is set to 100 for fine-tuning, meta-learning and 20 for in-context learning. The golden model uses the same overall number of runs ($L$) (1 000 for fine-tuning and meta-learning, 200 for in-context learning). These values, selected based on an Ablation Study (included in Appendix C), provide a balance between the coverage of the configurations’ state space and the computation costs.
Experimental Setup.
Flan-T5 Random Fixed Interactions
Golden Model 2.244 2.244 2.244
Label Select. (*) 2.517 (*) 2.594 (*) 2.128
Data Split (*) 2.362 (*) 2.480 (*) 2.167
Data Order (*) 2.131 (*) 3.014 0.869
Sample Choice (*) 2.370 (*) 3.191 (*) 2.123
Zephyr-7B Random Fixed Interactions
Golden Model 1.043 1.043 1.043
Label Select. (*) 1.122 (*) 1.004 (*) 0.863
Data Split (*) 1.185 0.402 (*) 0.664
Data Order (*) 1.138 (*) 0.957 0.456
Sample Choice (*) 1.052 0.406 (*) 0.744
Table 1: Comparison of different investigation strategies for the Flan-T5 and Zephyr-7B models on the SST2 dataset based on the F1 macro standard deviation. Factors considered important for different strategies are denoted using the (*) symbol. We observe that interactions between factors may cause some factors to have their importance overestimated (denoted in bold) or underestimated (denoted in italics).
4.1 Interactions Between Randomness Factors
In this section, our goal is to answer the following research question: RQ1: How do the interactions between randomness factors affect their individual importance? To answer this question, we compare our proposed method (Interactions) with the commonly used investigation strategies: 1) Random, which varies the overall random seed in the investigation without any constraint on the configurations of other factors; and 2) Fixed, where the non-investigated randomness factors are fixed to a single configuration for all runs of the investigation. For these strategies, we consider the effects of factor to be important when it contributes at least 50% of the golden model standard deviation. The results from this comparison are shown in Table 1 and used for validation of our method in Appendix E.
Effects of randomness factors may be overestimated or underestimated when interactions are not taken into consideration. The Random strategy leads to a deviation similar to the one from the golden model across all investigated randomness factors. Such result indicates that all randomness factors are equally important, leading to a significant importance overestimation in some cases (e.g., Data Order factor for both Flan-T5 and Zephyr models). Even though the Fixed strategy produces more reliable results, it is still affected by the random choice of the single factor configuration (e.g., Data Order contributing deviation of $3.014$, which is much higher than the deviation of $2.244$ from the golden model). As such, we observe underestimation of the results (e.g., Sample Choice and Data Split with the Zephyr-7B model not being considered important with a deviation of $0.406$ and $0.402$) as well as overestimation (e.g., Data Order being considered important with a deviation of $3.014$ for Flan-T5 and $0.957$ for Zephyr-7B). Taking the interactions into consideration, we observe that the Data Order randomness factor is not consistently important for in-context learning even when choosing samples randomly, which confirms the impact of interactions on incorrect attribution of effects of different randomness factors.
4.2 Importance of Randomness Factors
4.3 Effects of Variable Number of Classes and In-Context Samples
In this section, our focus is on answering the following research question: RQ3: How does the importance of data-specific randomness factors change based on the number of classes and in-context samples? As we observe different effects of randomness factors on binary and multi-class datasets, our main focus is to determine whether the change in importance is caused by the increased number of in-context examples in the prompt, by the larger number of options that can be predicted or by a combination of both. The results from changing the number of classes are included in Figure 1, and from changing the number of shots for in-context learning are included in Figure 2.
The importance of the Sample Choice for in-context learning is not consistently affected by the number of classes. In case of Flan-T5 and Zephyr, the importance of Sample Choice gradually decreases as we increase the number of classes (from $0.57$ on SST2 to $-0.57$ on SNIPS for Flan-T5, or from $0.39$ on SST2 to $0.17$ on DB-Pedia for Zephyr). For the LLaMA-2 model, the decrease is not as consistent, with the importance being much lower on the TREC than on the SNIPS dataset. Finally, the Sample Choice randomness factor is consistently important across all datasets for the Mistral model, with no apparent tendency.
4.4 Impact of Prompt Format
4.5 Discussion
Besides understanding the sensitivity, an important aspect is predicting a good configurations of the most important randomness factors to guarantee stability and generalisability of the approaches. As opposed to the hyperparameter tuning, finding this configuration is not as straightforward as it is not so systematic. First, as we show in this work, the importance and the best configuration is strongly affected by the interactions between randomness factors and other systematic choices. Second, there is no metric that can serve as an estimate for the quality of the configuration besides the observed performance – for example when finding an optimal prompt, the best performing and worst performing one can differ only in a single word (Zhan et al., 2024).
One way how to determine the optimal configuration are the mitigation strategies that recently started to attract a research attention (see the recent survey on stability of learning with limited labelled data (Pecher et al., 2024b) for more information). While the mitigation strategies are often factor-specific, such as sample selection strategies for in-context learning (Li and Qiu, 2023; Chang and Jia, 2023; Köksal et al., 2023; Pecher et al., 2024c), also more general strategies based on ensembling and further model training have been developed (Pecher et al., 2024a; Pezeshkpour and Hruschka, 2023; Summers and Dinneen, 2021; Wang et al., 2023; Allingham et al., 2023; Voronov et al., 2024). Uncovering the most important randomness factors through systematic investigation and design of new and effective mitigation strategies to reduce the sensitivity to the effects of randomness is an important future directions of the field (Pecher et al., 2024b; Liu et al., 2023).
5 Conclusion
In this work, we have proposed a novel method that explicitly takes interactions between different randomness factors into consideration by mitigating the effects of the other, non-investigated, randomness factors. In addition, our method is designed to determine the importance of the investigated randomness factor by measuring what fraction of the overall deviation of the model (represented using a golden model) it contributes over the mitigated randomness factors, allowing for in-depth analysis across experimental settings.
Applying our proposed method to investigate the effects of randomness factors on in-context learning, fine-tuning and meta-learning, we confirm our hypothesis that interactions between randomness factors may cause incorrect attribution of effects of one factor to another, leading to inconsistent result. Contrary to the previous works, after taking interactions into consideration, we do not observe a consistent sensitivity of in-context learning approaches to the sample order even when choosing samples at random. Instead, we observe that the importance of randomness factors, especially the sample order, is affected by the interactions with other factors and by the systematic choices such as the number of predicted classes, the number of samples per class and the choice of prompt format.
The proposed method can be applied to other NLP tasks as well, such as question answering, with minimal modifications. Only requirement is to define the randomness factors and their configurations, such as order of choices in the questions or the symbols used for the answers. Extending our investigation to other tasks represents an interesting potential for future work.
Acknowledgements
This work was partially supported by the projects funded by the European Union under the EU Horizon 2020: TAILOR, GA No. 952215, by the European Union under the Horizon Europe: DisAI, GA No. 101079164 and vera.ai, GA No. 101070093, and by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I03-03-V03-00020.
Limitations
The effects of randomness factors are investigated on a selection of models from different approaches for learning with limited labelled data. However, in order to provide a more extensive and in-depth analysis of the interactions and the more systematic choices without a significant increase in computation costs, the effects are investigated on models of smaller sizes – we use the base versions of BERT, RoBERTa and Flan-T5 models, and a 4-bit quantised versions of the LLaMA-2-13B, Mistral-7B and Zephyr-7B models. As such, the observed effects may not be as representative for larger models. However, similar to related work, we observed the larger models to be more susceptible to the effects of randomness and so the results of our investigation may underestimate the importance of different factors (instead of their over-estimation).
The number of investigation and mitigation runs used in our investigation is selected based on an ablation study (in Appendix 3). In addition, following related work (e.g., Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2023); Li and Qiu (2023); Köksal et al. (2023)) and the results of our ablation study, we also evaluate each run using only 1 000 test samples. In both cases, the decision represents a trade-off between the reliability of the results and their feasibility. Although this represents an optimal trade-off, increasing the number of runs and the number of test samples could potentially lead to better estimation of the effects and mitigation of their interactions (especially on larger datasets), but at the cost of significant computation costs increase. At the same time, the number of investigation and mitigation runs utilised in this paper still represents a significant improvement over the existing studies, as it is a common practice to investigate the effects using very low numbers of runs. As future work, we plan to explore the possibilities for effective mitigation strategies for all the randomness factors to mitigate their effects and further reduce computation costs.
Appendix A Ethical Considerations and Impact Statement
Impact Statement: CO2 Emissions Related to Experiments
The experiments presented in this paper used significant compute resources as they required multiple training and evaluation runs of multiple models, as well as using large language models that require a lot of computation even just for the inference. Overall, the experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh (default value used as the actual efficiency of our HW instance was not measured). A cumulative of 1440 hours of computation was performed on hardware of type RTX 3090 (TDP of 350W) and a cumulative of 4000 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W). The hours of computation used are only a crude approximation, as the machine used was shared among multiple projects. Total emissions are estimated to be 217.73 kgCO2eq (for the first set of hardware) and 432 kgCO2eq (for the second set of hardware), of which 0 percents were directly offset. These estimations were conducted using the MachineLearning Impact calculator presented in Lacoste et al. (2019). Whenever possible, we tried to reduce the compute resources used as much as possible. The most compute resources were used by the large language model – LLaMA-2, Mistral-7B and Zephyr-7B. To reduce the computation costs and resources used, we decided to evaluate the model on lower number of runs (10 investigation and 20 mitigation, resulting in 200 runs for each randomness factors) using only 1 000 test samples for evaluation. Even in this reduced evaluation, the experiments using these models used the most GPU hours. To further reduce the compute resources, we use the 4-bit quantised versions of these models, while also opting to use smaller models for the more detailed analyses and ablation studies, either in case of in-context learning (e.g., using Flan-T5 and Mistral-7B instead of LLaMA-2 for studying the impact of number of shots that significantly increased the required computation resources and inference time), but also in case of transfer learning and meta-learning (e.g., using base versions of the BERT and RoBERTa models).
Appendix B Additional Resources Describing the Proposed Investigation Method
In this Appendix, we provide additional supplementary resources that should allow for easier understanding of how the method operates. The proposed investigation method is designed for investigating effects of any randomness factor, while explicitly taking interactions with effects of other factors into consideration, and measuring the importance of the found effects. Overall, the effects are investigated by observing how the performance changes across the different states the investigated randomness factors can appear in. To deal with the interactions, the effects of remaining randomness factors are mitigated (i.e., the deviation they contribute is reduced as much as possible). To determine the importance, we compare the contributed deviation of the investigated randomness factor with effects of other factors, and with the overall deviation from a golden model. The golden model represents the objective estimate of the performance deviation and is obtained by training a model while mitigating all the randomness factors at the same time. The final importance score is then determined as the fraction of the overall deviation (represented using the golden model) the investigated factor contributes over all the remaining, non-investigated factors.
The following section provides a high-level overview of the method with references to the Algorithm 1 (Appendix B.1), the illustration of how the method operates and the results it computes (Appendix B.2). We also provide a method for selecting the number of investigation and mitigation runs in Appendix B.3 (this method was used to select the samples in this paper using the Ablation Study in Appendix 3 and to produce the heuristics at the end of Section 3).
B.1 Algorithmic Description of the Method
To allow for better understanding of our proposed investigation method, we provide more informal description of the steps composed in Algorithm 1, along with references to individual lines in it and possible avenues for extension of our method. Informally our proposed method works in a following way:
  1. 1.
    A set of randomness factors for investigation is first identified along with their configurations. In case of mitigated randomness factors, a complete set of factors and their configurations is not required to prevent introduction of biases into the results, as the randomness factors can be controlled on the group level. All the algorithmic factors (order, initialisation, model randomness, etc.) can be controlled by globally setting the seed, while the implementation/hardware level factors can be controlled using the same setup across all experiments (same library versions, architectures, GPUs, etc.).
  2. 2.
  3. 3.
    The method can be easily extended to investigate effect of multiple factors at the same time, by simply changing the definition of the investigated factor configuration set, to include a cross product between the configurations of multiple factors (similarly to the mitigated factor configuration set). (lines 2 and 3 in the Algorithm 1)
  4. 4.
  5. 5.
  6. 6.
    Instead of repeating multiple mitigation runs, the method can be extended to use a specific mitigation strategy (such as sample selection method for in-context learning). Using such method, the set of configurations for the given randomness factor is simply replaced by the results of the mitigation strategy (either a set of single value or a subset that is significantly smaller). The rest of the method remains unchanged. (line 3 in the Algorithm 1)
  7. 7.
  8. 8.
    The golden model standard deviation is calculated by simply performing training and evaluation multiple times with differently fixed configuration of all randomness factors. If enough overall runs are used, the golden model standard deviation can be replaced by simply taking the standard deviation over all the runs in the investigation. However, this may lead to incorrect results. (final loop; lines 13-17 in the Algorithm 1)
  9. 9.
@@ B.1 Algorithmic Description of the Method
    The importance score of the investigated factor is calculated as a fraction of the golden model standard deviation of the difference between contributed standard deviation and the mitigated standard deviation (to determine how much more the investigated factor contributes over all the mitigated ones). Any randomness factor with importance score over 00 is considered significantly important, as such factors contribute the same amount of deviation as the combination of all the remaining factors. At the same time, the size of the importance value determines the overall importance of the model (i.e., factor with importance score of $0.6$ is more important than the ones with score of $0.1$). (the final check; lines 18-21 in the Algorithm 1)
B.2 Illustration of the Method and its Results
@@
In this section, we provide the visualisation of the method in a form of table. In essence, when investigating the specific randomness factor, while mitigating the effects of other randomness factors, we fill in such table as illustrated in Table 2. The columns represent the different configurations for the investigated factor. Observing how the performance changes across these columns, we can determine the effects of the randomness factors – aggregating across these columns we obtain the partial mean $p\_mean$ and partial standard deviation $p\_std$.
@@
B.3 Selecting Number of Investigation and Mitigation Runs
When selecting the number of investigation ($N$) and the number of mitigation ($M$) runs, we need to find a balance between how well the effects of the factors are estimated and how well the interaction between the effects of different randomness factors are mitigated, and how much computational resources are required to get to this estimation and mitigation. An optimal solution is to use the lowest number of overall runs (that lead to lowest computational resources) after which the change in the results (the contributed/mitigated standard deviation or the normalised importance score) is under an acceptable threshold $\epsilon$. The value of this threshold $\epsilon$ depends on the setup of the experiment and the goal of our investigation, as in some cases higher change in the standard deviation may be acceptable, while in others we require a more strict setting.
In this section, we describe a simple method to search for this optimal point that can be used instead of the heuristics at the end of Section 3 (which were a result of our analysis using the following method). The method is composed of following steps:
  1. 1.
    The threshold of smallest acceptable change $\epsilon$, and the starting number of investigation runs $N$ are selected. The number of investigation runs should be sufficiently high from the start (following recommendations in Section 3) to make the search faster.
  2. 2.
    A new mitigation run should be performed using a randomly selected configuration of the non-investigated randomness (or the number of investigation runs should be increased, running the new investigation runs for all the already performed mitigation runs).
  3. 3.
    The new values of the relevant metrics (contributed standard deviation, mitigated standard deviation, or the importance score) should be determined and the difference to previous values calculated.
  4. 4.
    If the observed change is lower than the threshold $\epsilon$ the current values of hyperparameters $N$ and $M$ represent the optimal point and should be used. Otherwise, continue to step 2 (increasing either the value of $N$ or $M$).
In case our goal is to use the results of the investigation and the importance score for a more in-depth analysis and comparison across different factors, models, datasets or other experimental settings, the method should be repeated for every setting and the highest values of $N$ and $M$ should be used – to guarantee that the comparison and analysis is done on the same number of overall runs and to not introduce any possible biases into the comparison.